​(Standard deviation​) Given the following probabilities and returns for​ Mik's Corporation, find the standard deviation. 

  PROBABILITY 0.15 . 0.45 . 0.25 . 0.15

RETURNS .   10​% . 5​% . 18​% . 13​%

​Mik's standard deviation is ------%. ​(Round to two decimal​ places)
Answer #1

Expected Return = 0.15 * 0.10 + 0.45 * 0.05 + 0.25 * 0.18 + 0.15 * 0.13
Expected Return = 0.1020 or 10.20%

Variance = 0.15 * (0.10 - 0.1020)^2 + 0.45 * (0.05 - 0.1020)^2 + 0.25 * (0.18 - 0.1020)^2 + 0.15 * (0.13 - 0.1020)^2
Variance = 0.002856

Standard Deviation = (0.002856)^(1/2)
Standard Deviation = 0.0534 or 5.34%